# Rethinking Multimodal Learning from the Perspective of Mitigating Classification Ability Disproportion

### v1

@@ 4.1 Sustained Boosting
5:     Set assigning video classifier as true.
6:else if $s^{a}_{t}-\sigma s^{v}_{t}<\tau$ then
7:     Set assigning audio classifier as true.
8:end if
@@
10:     Minimize ${\mathcal{L}}_{\textit{ACA}}$ according to SGD;
11:     Update ${\boldsymbol{\omega}}_{t}$;
12:until Converge.
4.2 Adaptive Classifier Assignment
Thus far, we have defined a configurable classifier module and designed a sustained boosting in MML to enhance the classification performance of weak modality. However, recent studies [31] have shown that differences between modalities evolve dynamically due to imbalance issues in MML. This implies the need to design a strategy for enhancing classification ability that adapts to dynamic changes. Hence, we propose an adaptive classifier assignment strategy to adjust the number of the weak classifier. For simplicity, we redefine the modality classifiers as: $\Phi^{r}_{t}\triangleq\{\theta^{r},\Theta^{r}_{t}\},t\in\{1,\cdots,n^{r}\}$, where $n^{r}$ is the parameter to be updated.
Then, we utilize OGM [31] score to monitor the learning status. At $t$-th iteration, OGM score can be calculated by:
$\forall r\in\{a,v\},s^{r}_{t}=\frac{1}{N}\sum_{i=1}^{N}{\bf y}_{i}^{\top}\left[\sum_{j=1}^{n^{r}}{\bf p}^{r}_{ij}\right].$
OGM score reflects the classification ability of the models. Hence, if $s^{a}_{t}-\sigma s^{v}_{t}>\tau$, we assign a new configurable classifier for video modality at this iteration, where $\sigma\geq 1$ is the coefficient. $\tau$ is the dead zone for fault tolerance. Unless otherwise specified, the default is $\sigma=1.0$ and $\tau=0.01$. On the contrary, we also assign a new configurable classifier for audio modality if $s^{a}_{t}-\sigma s^{v}_{t}<\tau$.
To precisely refine the imbalance between modalities, we use the learning state at the $t$-th iteration to compute the weights for each modality’s classifier. More specifically, we utilize multiple layers DNN to learn the weights. Corresponding objective can be formed as:
${\mathcal{L}}_{\textit{ACA}}({\bf s}_{t})=\|{\bf g}_{t}-{\boldsymbol{\omega}}_{t}\|^{2}_{2},$
where ${\boldsymbol{\omega}}_{t}=[\omega^{a}_{t},\omega^{v}_{t}]$ denotes the weight to be learned. And ${\bf g}_{t}=[g_{t}^{a},g_{t}^{v}]$ indicates the classification ability based on $s^{r}_{t}$, where $g_{t}^{a}=\delta(s_{t}^{a}=\mathtt{min}([s_{t}^{a},s_{t}^{v}]))$, $g_{t}^{v}=1-g_{t}^{a}$.
Then the prediction of each classifier is redefined as $\hat{\bf p}^{r}_{ij}=\omega^{r}_{t}{\bf p}^{r}_{ij}$. The learning algorithm of adaptive classifier assignment is summarized in Algorithm (1). Based on ACA algorithm, we can adjust the number of classifiers based on learning status and impose the weight over training of each modality by substituting ${\bf p}^{r}_{ij}$ as $\hat{\bf p}^{r}_{ij}$ in problem (6). Our algorithm is summarized in Algorithm (2). In practice, we perform ACA algorithm to determine if we need to adjust the classification ability every $T$ iterations.
Algorithm 2 Learning algorithm of our proposed method.
1:Training data ${\bf X}$, category labels ${\bf Y}$.
2:The learned DNN models for all modalities.
3:INIT initialize the number of classifier $n^{a}=1$, $n^{v}=1$. Initialize iteration $t=1$. Initialize DNN parameters $\Phi^{a}_{t}$ and $\Phi^{v}_{t}$. Initialize ${\boldsymbol{\omega}}_{t}=[1,1]$.
4:repeat
5:     // Learn MML models.
6:     Sample a mini-batch ${\bf X}_{t}=\{({\bf x}_{i}^{a},{\bf x}_{i}^{v})\}_{i=1}^{n_{b}}$;
7:     $\forall{\bf x}_{i}^{a},{\bf x}_{i}^{v}\in{\bf X}_{t}$, calculate features ${\bf u}_{i}^{a}$ and ${\bf u}_{i}^{v}$;
8:     Calculate predictions $\{{\bf p}^{a}_{ij}\}_{j=1}^{n^{a}}$ and $\{{\bf p}^{v}_{ij}\}_{j=1}^{n^{v}}$.
9:     Calculate loss in (6) based on predictions and ${\boldsymbol{\omega}}_{t}$;
10:     Update DNN parameters $\Phi^{a}_{t}$ and $\Phi^{v}_{t}$ based on SGD;
11:     // Call ACA algorithm.
12:     if $\mathtt{mod}(t,T)=0$ then
13:         Calculate OGM score based on predictions;
14:         Call ACA algorithm to update ${\boldsymbol{\omega}}_{t}$ and assignment decisions;
15:         if Assigning audio classifier is true then
16:              Add a classifier for audio modality;
17:              $n^{a}=n^{a}+1$;
18:         end if
19:         if Assigning video classifier is true then
20:              Add a classifier for video modality;
21:              $n^{v}=n^{v}+1$;
22:         end if
23:     end if
24:     Update $t=t+1$;
25:until Converge or reach maximum iterations.
4.3 Model Inference
After training, the learned multimodal models can be applied to perform classification for any unseen data point. More specifically, given data point ${\bf x}_{i}=({\bf x}_{i}^{a},{\bf x}_{i}^{v})$, we utilize the following equation to obtain the predictions:
$\forall r\in\{a,v\},\;\bar{\bf p}_{i}^{r}=\sum_{t=1}^{n^{r}}{\bf p}^{r}_{it}=\sum_{t=1}^{n^{r}}\psi^{r}_{t}(\phi^{r}({\bf x}_{i}^{r};\theta^{r});\Theta^{r}).$
Based on $\bar{\bf p}_{i}^{a}$, $\bar{\bf p}_{i}^{v}$, and the learned weights ${\boldsymbol{\omega}}^{*}$, we can adopt a specific late fusion strategy to obtain the final prediction.
5 Experiments
Table 1: The accuracy results on CREMAD, KSounds, and NVGesture datasets. The best accuracy is shown in boldface. The Top-3 results are highlighted with progressively darker shades of orange.
Method CREMAD KSounds NVGesture
Multi Audio Video Multi Audio Video Multi RGB OF Depth
G-Blend  [CVPR’20] 0.6465 0.6075 0.4301 0.6710 0.5160 0.4275 0.8299 0.7054 0.7178 0.7252
MSLR  [ACL’22] 0.6868 0.6357 0.2903 0.6756 0.5199 0.3254 0.8237 0.3672 0.3755 0.5373
OGM  [CVPR’22] 0.6612 0.6209 0.2903 0.6582 0.5013 0.3165 -- -- -- --
PMR  [CVPR’23] 0.6659 0.6263 0.4355 0.6675 0.4750 0.3772 -- -- -- --
AGM  [ICCV’23] 0.6733 0.4798 0.3655 0.6787 0.5036 0.3869 0.8279 0.6598 0.6722 0.7324
MMPareto  [ICML’24] 0.7487 0.6586 0.5108 0.7000 0.5226 0.4953 0.8382 0.7593 0.7925 0.8050
MLA  [CVPR’24] 0.7943 0.5727 0.6491 0.7004 0.5572 0.5402 0.8340 0.7241 0.7573 0.7742
ReconBoost  [ICML’24] 0.7557 0.5966 0.6364 0.6855 0.4941 0.5031 0.8386 0.7290 0.7471 0.7782
Ours 0.8441 0.6788 0.6770 0.7324 0.5219 0.5856 0.8575 0.7732 0.8070 0.8001
Table 2: The results on image-text datasets, i.e., Twitter and Sarcasm. The results are indicated similarly to those in Table 1.
Method Accuracy MacroF1
Twitter Sarcasm Twitter Sarcasm
G-Blend 0.7309 0.8286 0.6799 0.8215
MSLR 0.7232 0.8439 0.6382 0.8378
OGM 0.7058 0.8360 0.6435 0.8293
PMR 0.7357 0.8310 0.6636 0.8256
AGM 0.7261 0.8360 0.6502 0.8293
MMPareto 0.7358 0.8348 0.6729 0.8284
MLA 0.7352 0.8426 0.6713 0.8348
ReconBoost 0.7442 0.8437 0.6832 0.8317
Ours 0.7450 0.8450 0.6794 0.8384
5.1 Dataset
We carry out the experiments on six extensive multimodal datasets, i.e., CREMAD [4], KSounds [2], NVGesture [29], VGGSound [6], Twitter [44], and Sarcasm [3] datasets. The CREMAD, KSounds, and VGGSound datasets consist of audio and video modalities. NVGesture dataset contains three modalities, i.e., RGB, optical flow (OF), and Depth. Twitter and Sarcasm datasets consist of image and text modalities. The CREMAD dataset contains 7,442 clips, which are divided into training set with 6,698 samples and testing set with 744 samples. For KSounds dataset, which contains 19,000 video clips, is divided into training set with 15,000 clips, validation set with 1,900 clips, and testing set with 1,900 clips. VGGSound dataset includes 168,618 videos for training and validation, and 13,954 videos for testing. The NVGesture dataset is divided into 1,050 samples for training and 482 samples for testing.Twitter dataset is divided into training set with 3,197 pairs, validation set with 1,122 pairs and testing set with 1,037 pairs. Sarcasm dataset includes 19,816 pairs for the training set, 2,410 pairs for the validation set, and 2,409 pairs for the testing set. More details are provided in the appendix.
5.2 Experimental Settings
5.2.1 Baselines and Evaluation Metric
We select various SoTA baselines for comparison, including G-Blend [37], MSLR [43], OGM [31], PMR [11], AGM [26], MMPareto [38], MLA [45], and ReconBoost [19]. Among these methods, ReconBoost employs the gradient boosting algorithm to capture the error caused by other modalities.
Following the setting of MLA [45] and ReconBoost [19], we adopt accuracy, mean average precision (MAP) and MacroF1 as evaluation metrics. The accuracy measures the proportion of correct predictions of total predictions. MAP returns the average precision of all samples. And MacroF1 calculates the average F1 across all categories.
5.2.2 Implementation Details
Following OGM [31], we employ ResNet18 [17] as the backbone to encode audio and video for CREMAD, KSounds and VGGSound datasets. All the parameters of the backbone are randomly initialized. For NVGesture dataset, we employ the I3D [5] as unimodal branch following the setting of [40]. We initialize the encoder with the pre-trained model trained on ImageNet. For the architecture of the configurable classifier, we explore a two-layer network, which can be denoted as “Layer1$(D\times 256)\mapsto$ ReLU $\mapsto$ Layer2 $(256\times K)$”. Here, $D$ denotes the output dimensions of encoders, “Layer1”/“Layer2” are fully connected layer, and “ReLU” denotes the ReLU [1] activation layer. Furthermore, the Layer2 is utilized as shared head for all modalities as described in Section 4. Both Layer1 and Layer2 are randomly initialized. In addition, all hyper-parameters are selected by using the cross-validation strategy. Specifically, we use stochastic gradient descent (SGD) as the optimizer with a momentum of $0.9$ and weight decay of $1\times 10^{-4}$. The initial learning rate is set to be $1\times 10^{-2}$ for CREMAD, KSounds, VGGSound , and NVGesture datasets. During training, the learning rate is progressively reduced by a factor of ten upon observing loss saturates. The batch size is set to be 64 for CREMAD and KSounds datasets, 16 for VGGSound dataset, and 2 for NVGesture dataset. We set the iteration $T$ for checking whether to assign the classifier to 20 epochs for CREMAD dataset, 10 for KSounds, NVGesture datasets. For all datasets, we set $\lambda$ to be $0.2$. For the ACA algorithm in (1), we adopt a three-layer network with ReLU to learn ${\boldsymbol{\omega}}_{t}$, and the init learning rate is $0.001$. The optimization algorithm of ACA is the same as that of the backbone. For Twitter and Sarcasm datasets, following [44, 3], we adopt BERT [10] as the text encoder and ResNet50 [17] as the image encoder. We use Adam [25] as the optimizer, with an initial learning rate of $1\times 10^{-5}$. The batch size is set to 64. We set iteration $T$ as 5 for Twitter and 1 for Sarcasm dataset. The other parameter settings are the same as audio-video datasets. For comparison methods, the source codes of all baselines are kindly provided by their authors. For fair comparison, all baselines also adopt the same backbone and initialization strategy for the experiment. All experiments are conducted on an NVIDIA GeForce RTX 3090 and all models are implemented with pytorch.
Table 3: The accuracy on VGGSound dataset. The accuracy is indicated similarly to those in Table 1.
Method Accuracy
Multi Audio Video
AGM 0.4711 0.4548 0.2344
MLA 0.5165 0.4675 0.2616
MMPareto 0.5125 0.4735 0.2485
ReconBoost 0.5097 0.4535 0.2263
Ours 0.5304 0.4747 0.2515
5.3 Experimental Results
Classification Performance Comparison: The classification results on all datasets are reported in Table 1, Table 2, and Table 3, where the best result is denoted as boldface, and the top-3 results are highlighted with progressively darker shades of orange. In Table 1, “--” denotes that corresponding methods cannot applied to the dataset with more than two modalities.
The results in Table 1 show the unimodal and multimodal accuracy on CREMAD, KSounds, and NVGesture datasets. From Table 1, we can see that: (1). Our method can outperform existing SoTA baselines to achieve the best accuracy in all cases for multimodal situations. Specifically, compared with the best baseline, our method achieves absolute boosts of 5.38%, 3.8%, and 2.86% on three datasets respectively; (2). Our method can achieve the best accuracy for unimodal situations in almost all cases except audio on KSounds dataset; (3). The accuracy on NVGesture dataset demonstrates that our method can extend to the case with more than two modalities and achieve the best performance.
We further present the multimodal classification performance on image-text datasets, i.e., Twitter and Sarcasm datasets, in Table 2. The results in Table 2 demonstrate that: (1). Compared with SoTA baselines, our method can achieve the best performance in terms of accuracy and MacroF1 in all cases for datasets with image and text modalities; (2). The absolute improvement on Twitter and Sarcasm datasets is relatively smaller than that on CREMAD and KSounds datasets. One possible reason behind this phenomenon is that the modality imbalance between audio and video modalities is more serious than that between image and text modalities.
As VGGSound is a relatively large dataset, we only choose a set of recent algorithms, including AGM [26], MLA, MMPareto, ReconBoost, for experiments. The results are shown in Table 3. We can find that compared with SoTA baselines, our method can achieve the best performance, demonstrating the effectiveness of our method in the case of large-scale datasets.
More results with different metrics and error bar are provided in the appendix due to space limitations.
5.4 Sensitivity to Hyper-Parameters
Sensitivity to Threshold $\sigma$: We study the influence of threshold $\sigma$ on CREMAD dataset. The accuracy with different $\sigma\in[1,1.75]$ is shown in Figure 3 (a). We can find that our method is not sensitive to threshold $\sigma$ in a large range.
(a). Threshold $\sigma$.
(b). Smoothing factor $\lambda$.
Figure 3: Sensitivity to hyper-parameter $\sigma$ and $\lambda$.
Table 4: The results for ablation study on CREMAD dataset.
Metric $\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ Multi Audio Video
0.8240 0.6721 0.6842
0.8246 0.6795 0.6714
Accuracy 0.8320 0.6794 0.6761
0.8300 0.6754 0.6781
0.8441 0.6788 0.6770
0.9006 0.7247 0.7658
0.8947 0.7382 0.7568
MAP 0.9066 0.7378 0.7559
0.8994 0.7365 0.7625
0.9121 0.7501 0.7547
Sensitivity to Smoothing Factor $\lambda$: We explore the influence of smoothing factor $\lambda$ on CREMAD dataset. The accuracy with different $\lambda\in[0.1,1]$ is reported in Figure 3 (b). We can find that our method is not sensitive to hyper-parameter smoothing factor $\lambda$ in a large range.

### v2

@@ 3.3 Adaptive Classifier Assignment
11:   Calculate confident score $\{s^{a}_{t},s^{v}_{t}\}$ based on predictions;
12:   if $s^{a}_{t}-\sigma s^{v}_{t}>\tau$ then
13:     Add a classifier for audio modality;
14:     $n^{a}=n^{a}+1$;
15:   else if $s^{a}_{t}-\sigma s^{v}_{t}<\tau$ then
16:     Add a classifier for video modality;
17:     $n^{v}=n^{v}+1$;
18:   end if
19:  end if
20:end for
3.4 Theoretical Analysis
The sustained boosting algorithm is introduced to reduce the loss gap between different modalities. In this section, we theoretically analyze its effect on minimizing this gap. We first define the gap function as follows:
${\mathcal{G}}(\Phi)={\mathcal{L}}^{a}(\Phi^{a})-{\mathcal{L}}^{v}(\Phi^{v}),$ (7)
where $\Phi^{a}$ and $\Phi^{v}$ respectively denote the parameters of audio and video modality, ${\mathcal{L}}^{a}$ and ${\mathcal{L}}^{v}$ denote the overall loss function ${\mathcal{L}}_{\mathrm{SUB}}$ for audio and video, respectively.Without loss of generality, we assume that ${\mathcal{L}}^{a}>{\mathcal{L}}^{v}$; the case where ${\mathcal{L}}^{a}<{\mathcal{L}}^{v}$ can be analyzed analogously.
We derive the following conclusions:
Theorem 1 (Convergence of Gap Loss, Informal)
Under some assumptions for the loss function and the effectiveness of sustained boosting algorithm, we have:
${\mathcal{G}}(\Phi(T))\leq\frac{1}{1+\frac{\nu^{2}\kappa^{2}}{2L_{a}\beta^{2}}T{\mathcal{G}}(\Phi(0))}{\mathcal{G}}(\Phi(0))$ (8)
where $\nu$, $\kappa$, $L_{a}$ and $\beta$ are constant.
The results in Theorem 1 indicate that, by employing the gradient boosting algorithm, the loss gap between modalities converges at a rate of ${\mathcal{O}}(1/T)$. The proof can be found in the appendix.
4 Experiments
4.1 Experimental Setup
Dataset: We carry out the experiments on six extensive multimodal datasets, i.e., CREMAD CREMAD:journals/taffco/CaoCKGNV14 , KSounds Kinetics-Sound:conf/iccv/ArandjelovicZ17 , NVGesture NVGeasture:conf/cvpr/MolchanovYGKTK16 , VGGSound VGGSound:conf/icassp/ChenXVZ20 , Twitter Twitter15:conf/ijcai/Yu019 , and Sarcasm Sarcasm:conf/acl/CaiCW19  datasets. The CREMAD, KSounds, and VGGSound datasets consist of audio and video modalities. NVGesture dataset contains three modalities, i.e., RGB, optical flow (OF), and Depth. Twitter and Sarcasm datasets consist of image and text modalities.
The CREMAD dataset contains 7,442 clips, which are divided into training set with 6,698 samples and testing set with 744 samples. For KSounds dataset, which contains 19,000 video clips, is divided into training set with 15,000 clips, validation set with 1,900 clips, and testing set with 1,900 clips. VGGSound dataset includes 168,618 videos for training and validation, and 13,954 videos for testing. The NVGesture dataset is divided into 1,050 samples for training and 482 samples for testing. Twitter dataset is divided into training set with 3,197 pairs, validation set with 1,122 pairs and testing set with 1,037 pairs. Sarcasm dataset includes 19,816 pairs for the training set, 2,410 pairs for the validation set, and 2,409 pairs for the testing set. More details are provided in the appendix.
Baselines: We select two categories of methods for comparison, i.e., traditional multimodal fusion methods and rebalanced multimodal learning methods. In detail, traditional multimodal fusion methods include concatenation fusion (Concat), affine transformation fusion (Affine) Film:conf/aaai/PerezSVDC18 , multi-layers lstm fusion (ML-LSTM) ML-LSTM:journals/mta/NieYSW21 , prediction summation fusion (Sum) MML:conf/ijcai/YangWZX019 , and prediction weighting fusion (Weight) MML:conf/ijcai/YangWZX019 . Rebalanced multimodal learning methods include MSES MSES:conf/acpr/FujimoriEKM19 , G-blend OGR-GB:conf/cvpr/WangTF20 , MSLR MSLR:conf/acl/YaoM22 , OGM OGM:conf/cvpr/PengWD0H22 , PMR PMR:conf/cvpr/Fan0WW023 , AGM AGM:conf/iccv/LiLHLLZ23 , MMParato MMPareto:conf/icml/WeiH24 , SMV SMV:conf/cvpr/YakeRZD24 , MLA MLA:conf/cvpr/ZhangYBY24 , DI-MML DI-MML:conf/mm/FanXWLG24 , LFM LFM:conf/nips/0074WJ024 , ReconBoost ReconBoost:conf/icml/CongHua24 .
Evaluation Protocols: Following the setting of MLA MLA:conf/cvpr/ZhangYBY24 and ReconBoost ReconBoost:conf/icml/CongHua24 , we adopt accuracy, mean average precision (MAP) and MacroF1 as evaluation metrics. The accuracy measures the proportion of correct predictions of total predictions. MAP returns the average precision of all samples. And MacroF1 calculates the average F1 across all categories.
@@ 4.1 Experimental Setup
Table 1: Classification performance comparison with SOTA baselines. The best and second-best results are highlighted in bold and underline, respectively. The results with gray background denote the performance based on multimodal learning is inferior to that of the best unimodal approach.
Method CREMAD KSounds VGGSound Twitter Sarcasm NVGesture
Acc. MAP Acc. MAP Acc. MAP Acc. F1 Acc. F1 Acc. F1
Unimodal-1 .4583 .5879 .5412 .5669 .4655 .4701 .5863 .4333 .7181 .7073 .7822 .7833
Unimodal-2 .6317 .6861 .5562 .5837 .3494 .3478 .7367 .6849 .8136 .8056 .7863 .7865
Unimodal-3 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A .8154 .8183
Concat .6361 .6841 .6455 .7130 .5116 .5352 .7011 .6386 .8286 .8240 .8237 .8270
Affine .6626 .7193 .6424 .6931 .5001 .5155 .7203 .5992 .8240 .8188 .8278 .8281
ML-LSTM .6290 .6473 .6394 .6902 .4966 .5139 .7068 .6564 .8277 .8205 .8320 .8330
Sum .6344 .6908 .6490 .7103 .5136 .5338 .7300 .6661 .8294 .8247 .8050 .8067
Weight .6653 .7134 .6533 .7110 .5144 .5300 .7242 .6516 .8265 .8219 .7842 .7939
MSES .6546 .7138 .6591 .7196 .4891 .5429 .7252 .6439 .8423 .8369 .8112 .8147
G-blend .6465 .7392 .6722 .7274 .5086 .5555 .7309 .6799 .8286 .8215 .8299 .8305
MSLR .6868 .7412 .6756 .7282 .4987 .5415 .7232 .6382 .8439 .8378 .8237 .8284
OGM .6612 .7372 .6582 .7159 .4829 .4978 .7058 .6435 .8360 .8293 - -
PMR .6659 .7058 .6675 .7274 .4647 .4866 .7357 .6636 .8310 .8256 - -
AGM .6733 .7807 .6791 .7388 .4711 .5198 .7261 .6502 .8306 .8293 .8279 .8284
MMParato .7487 .8535 .7000 .7850 .5125 .5473 .7358 .6729 .8348 .8284 .8382 .8424
SMV .7872 .8417 .6900 .7426 .5031 .5362 .7428 .6817 .8418 .8368 .8352 .8341
MLA .7943 .8572 .7004 .7945 .5165 .5473 .7352 .6713 .8426 .8348 .8340 .8372
DI-MML .8158 .8592 .7203 .7426 .5173 .5479 .7248 .6686 .8411 .8315 - -
LFM .8362 .9006 .7253 .7897 .5274 .5694 .7501 .7057 .8497 .8457 .8436 .8468
ReconBoost .7557 .8140 .6855 .7662 .5097 .5387 .7442 .6832 .8437 .8317 .8386 .8434
Ours
.8515
$\pm$0.0027
.9103
$\pm$0.0014
.7263
$\pm$0.0031
.7901
$\pm$0.0063
.5301
$\pm$0.0006
.5826
$\pm$0.0017
.7512
$\pm$0.0068
.6962
$\pm$0.0016
.8510
$\pm$0.0054
.8458
$\pm$0.0047
.8501
$\pm$0.0025
.8533
$\pm$0.0031
4.2 Main Results
Classification Performance Comparison: The classification results on all datasets are reported in Table 1, where “-” denotes that corresponding methods cannot applied to the dataset with more than two modalities. And Unimodal-1/2/3 is used to denote the results based on unimodal. Unimodal-1/2 respectively denote the video/audio for CREMAD and KSounds, and text/image for Twitter and Sarcasm. Unimodal-1/2/3 denotes the RGB/OF/Depth modality for NVGesture dataset, respectively. Furthermore, the results with a gray background indicate that the performance based on multimodal learning is inferior to that of the best unimodal approach. We can draw the following observations: (1). Compared with unimodal baselines, traditional multimodal fusion methods and rebalanced multimodal learning methods can achieve better performance in almost all cases; (2). Our method can outperform existing SOTA baselines to achieve the best accuracy in all cases for multimodal situations; (2). The accuracy on NVGesture dataset demonstrates that our method can extend to the case with more than two modalities and achieve the best performance.
4.3 Sensitivity to Hyperparameter
Sensitivity to Threshold $\sigma$: We study the influence of threshold $\sigma$ on CREMAD dataset. The accuracy with different $\sigma\in[1,1.75]$ is shown in Figure 3. We can find that our method is not sensitive to threshold $\sigma$ in a large range.
Sensitivity to Smoothing Factor $\lambda$: We explore the influence of smoothing factor $\lambda$ on CREMAD dataset. The accuracy with different $\lambda\in[0.1,1]$ is reported in Figure 3. We can find that our method is not sensitive to hyper-parameter smoothing factor $\lambda$ in a large range.
Figure 2: Sensitivity to $\sigma$ (left) and $\lambda$ (right).
Figure 3: Performance comparison.